# Exact Thresholds for Noisy Non-Adaptive Group Testing

### 10 Jan 2024 v1

@@ Appendix D $-\ell$ Achievability Analysis for Bernoulli Designs
Bounding the overall failure probability for $\ell\in[1,\frac{k}{\log k}]$: We further take a union bound over $\ell\in[1,\frac{k}{\log k}]$. We assume that the condition (152) holds for all $\ell\in[1,\frac{k}{\log k}]$, i.e.,
$n\geq(1+\eta_{1})\frac{k}{\nu e^{-\nu}}\max_{1\leq\ell\leq\frac{k}{\log k}}\max_{\begin{subarray}{c}0<C<C_{0}\\ 0<\zeta<1\end{subarray}}\min\Big{\{}\frac{\log\frac{k}{\ell}}{f_{1}(C,\zeta,\rho)},\frac{\log\frac{p}{\ell}}{\hat{f}_{2}(C,\zeta,\rho)}\Big{\}}$ (157)
$\iff$ $n\geq\frac{(1+\eta_{1})k\log\frac{p}{k}}{(1-\theta)\nu e^{-\nu}}\frac{1}{\min_{C\in(0,C_{0}),\zeta\in(0,1)}\max\{\frac{1}{\theta}f_{1}(C,\zeta,\rho),\hat{f}_{2}(C,\zeta,\rho)\}},$ (158)
@@
where in (158) we observe that the maximum over $\ell$ is attained at $\ell=1$ and use $\log p\sim\frac{1}{1-\theta}\log\frac{p}{k},\leavevmode\nobreak\ \log k\sim\frac{\theta}{1-\theta}\log\frac{p}{k}$. Then, if (158) holds, by union bound we can bound the probability that the restricted MLE decoder fails as
$\mathbb{P}({\rm err})$ $\leq\mathbb{P}(\mathscr{A}_{1}^{c})+\sum_{\ell=1}^{k/\log k}\mathbb{P}\Big{(}\big{\{}\big{|}s\setminus\widehat{S}^{\prime}\big{|}=\ell\big{\}}\cap\mathscr{A}_{1}\Big{)}$ (159)
$\leq o(1)+\sum_{\ell=1}^{k/\log k}\Big{(}k^{-5}+O((\ell\log k)^{2})\hat{P}_{\ell}\Big{)}$ (160)
$\leq o(1)+\Big{[}\sum_{\ell=1}^{\log k}(\ell\log k)^{-3}+\sum_{\ell=\log k}^{k/\log k}k^{-3}\Big{]}=o(1),$ (161)
@@
where (159) is due to a union bound, and we use (154) in (160) and then substitute $\hat{P}_{\ell}$ (given by (149)) in (161). In conclusion, if (158) holds for some $\eta>0$, then the restricted MLE decoder fails to recover $s=[k]$ with $o(1)$ probability.
@@
Simplifying $\hat{f}_{2}(C,\zeta,\rho)$ to $f_{2}(C,\zeta,\rho)$: Recall that $\hat{f}_{2}(C,\zeta,\rho)$ is given in (143). The aim of this step is to show that the minimum over $(C,\zeta)$ in (158) is not attained in the domain of $(1-2\zeta)C<2\rho-1$, and thus we can safely simplify the $\hat{f}_{2}(C,\zeta,\rho)$ to $f_{2}(C,\zeta,\rho)$ in the threshold (158). Consider $(C,\zeta)$ satisfying $(1-2\zeta)C\leq 2\rho-1$, i.e.,
$\zeta\geq\zeta^{\prime}:=\frac{1}{2}+\frac{1-2\rho}{2C}\geq\frac{1}{2},$ (162)
@@
in which we have $\hat{f}_{2}(C,\zeta,\rho)=0$ (see (143) and recall that $\hat{f}_{2}(C,\zeta,\rho)$ is continuous), and that $f_{1}(C,\zeta,\rho)$ is monotonically increasing with regard to $\zeta$ in $[\frac{1}{2},1)$ (see (96)). Therefore, we can compare the values of $f_{1}(C,\zeta,\rho)$ and $f_{2}(C,\zeta,\rho)$ over $(C,\zeta)$ and $(C,\zeta^{\prime})$ as
$f_{1}(C,\zeta,\rho)\geq f_{1}(C,\zeta^{\prime},\rho),\leavevmode\nobreak\ \hat{f}_{2}(C,\zeta,\rho)=\hat{f}_{2}(C,\zeta^{\prime},\rho)=0,$ (163)
@@
$\max\Big{\{}\frac{1}{\theta}f_{1}(C,\zeta,\rho),\hat{f}_{2}(C,\zeta,\rho)\Big{\}}\geq\max\Big{\{}\frac{1}{\theta}f_{1}(C,\zeta^{\prime},\rho),\hat{f}_{2}(C,\zeta^{\prime},\rho)\Big{\}}.$ (164)
@@
$\min_{C\in(0,C_{0}),\zeta\in(0,1)}\max\Big{\{}\frac{1}{\theta}f_{1}(C,\zeta,\rho),\hat{f}_{2}(C,\zeta,\rho)\Big{\}}$ (165)
=\,= $\min_{\begin{subarray}{c}C\in(0,C_{0}),\zeta\in(0,1)\\ \zeta\leq\frac{1}{2}+\frac{1-2\rho}{2C}\end{subarray}}\max\Big{\{}\frac{1}{\theta}f_{1}(C,\zeta,\rho),\hat{f}_{2}(C,\zeta,\rho)\Big{\}}$ (168)
=\,= $\min_{\begin{subarray}{c}C\in(0,C_{0}),\zeta\in(0,1)\\ (1-2\zeta)C\geq 2\rho-1\end{subarray}}\max\Big{\{}\frac{1}{\theta}f_{1}(C,\zeta,\rho),f_{2}(C,\zeta,\rho)\Big{\}}$ (171)
@@
Therefore, to ensure (158), it suffices to have
$n\geq\frac{(1+\eta)k\log\frac{p}{k}}{(1-\theta)\nu e^{-\nu}}\frac{1}{\min_{C>0,\zeta\in(0,1)}\max\{\frac{1}{\theta}f_{1}(C,\zeta,\rho),f_{2}(C,\zeta,\rho)\}}$ (175)
for some $\eta>0$. This establishes the second term in (4).
@@ Appendix E $-\ell$ Converse Analysis for Near-Constant Weight Designs
In near-constant weight design, each item is independently and uniformly placed at $\Delta=\frac{\nu n}{k}$ tests with replacement. As with the Bernoulli design, it suffices to analyze the MLE decoder (29) that finds $\widehat{S}$ having the most correct test outcomes (i.e., outcomes that would be obtained if there were no noise). By the symmetry of the test design with respect to re-ordering items, we can again consider $s=[k]$ as the underlying defective set without loss of generality.
@@
For $j\in s$ and $j^{\prime}\in[p]\setminus s$ we will use the notation $\mathcal{M}_{j},\mathcal{M}_{j0},\mathcal{M}_{j1},\mathcal{N}_{0},\mathcal{N}_{00},\mathcal{N}_{01}$ and the corresponding cardinalities (e.g., $N_{0}=|\mathcal{N}_{0}|$, $M_{j}=|\mathcal{M}_{j}|$) from Corollary 1. These index sets and their cardinalities are known under given $\mathbf{X}_{s}$ and $\mathbf{Y}$. We consider $\ell=1$, and say that $(C,\zeta)\in(0,\infty)\times(0,1)$ is feasible if $\frac{Cn\nu e^{-\nu}}{k}$ and $\frac{\zeta\cdot Cn\nu e^{-\nu}}{k}$ are integers. Given feasible $(C,\zeta)$ we will consider $\mathcal{K}_{C,\zeta}=\{j\in s:M_{j}=\frac{Cn\nu e^{-\nu}}{k},\leavevmode\nobreak\ M_{j0}=\frac{\zeta\cdot Cn\nu e^{-\nu}}{k}\}$ with cardinality $k_{C,\zeta}=|\mathcal{K}_{C,\zeta}|$ as in (35). For given $(j,j^{\prime})\in\mathcal{K}_{C,\zeta}\times([p]\setminus s)$, recall that $G_{j,j^{\prime},1}$ denotes the number of tests in $\mathcal{N}_{01}\cup\mathcal{M}_{j1}$ that contain the non-defective item $j^{\prime}$, and $G_{j,j^{\prime},2}$ denotes number of tests in $\mathcal{N}_{00}\cup\mathcal{M}_{j0}$ that contain the non-defective item $j^{\prime}$; see Corollary 1.
We need some additional conventions for analyzing the near-constant design. For $j\in s$ and $j^{\prime}\in[p]\setminus s$, we let $\mathcal{M}_{j,j^{\prime}}$ index the tests in $\mathcal{M}_{j}$ that contain item $j^{\prime}$, and $\mathcal{N}_{0,j^{\prime}}$ index the tests in $\mathcal{N}_{0}$ that contain item $j^{\prime}$. Then we note the equivalent interpretations of $G_{j,j^{\prime},1}$ and $G_{j,j^{\prime},2}$: $G_{j,j^{\prime},1}$ represents the number of positive tests in $\mathcal{N}_{0,j^{\prime}}\cup\mathcal{M}_{j,j^{\prime}}$, and $G_{j,j^{\prime},2}$ represents the number of negative tests in $\mathcal{N}_{0,j^{\prime}}\cup\mathcal{M}_{j,j^{\prime}}$; this alternative formulation happens to be more convenient for analyzing the near-constant weight design.
In addition, for $j\in s$, we let $\mathcal{M}_{j}^{\prime}$ index the tests where item $j$ not only appears as the only defective but is also placed precisely once, and let its cardinality be $M_{j}^{\prime}=|\mathcal{M}_{j}^{\prime}|$. Compared to Bernoulli designs, this is a new ingredient for handling near-constant weight designs whose random placement is done with replacement. Evidently, it holds that $M_{j}^{\prime}\leq M_{j}$, and $M_{j}^{\prime}<M_{j}$ occurs when item $j$ is placed in some test more than once but is still the only defective. This situation only occurs for a small number of defective items, as we will formalize via event $\mathscr{A}_{1}$ below.
@@
Similarly to the converse analysis for Bernoulli design, by Corollary 1, MLE fails if for some feasible $(C,\zeta)\in(0,e^{\nu})\times(0,1)$, the following two conditions (C1) and (C2) simultaneously hold:
@@
  • (C2) There exist $j\in\mathcal{K}_{C,\zeta}$ and $k+1\leq j^{\prime}\leq p$ such that
    $G_{j,j^{\prime},1}-G_{j,j^{\prime},2}>(1-2\zeta)\frac{Cn\nu e^{-\nu}}{k}.$ (176)
@@
Before proceeding, we first construct three high-probability events. By Lemma 5(a) in Appendix G, for some sufficiently large constant $C_{0}>0$, the event
$\mathscr{A}_{1}=\Big{\{}|\{j\in[k]:M_{j}^{\prime}<M_{j}\}|\leq C_{0}\log k\Big{\}}.$ (177)
@@
$\widetilde{M}:=\sum_{j=1}^{k}M_{j}^{\prime},$ (178)
which represents the number of tests connected to precisely one defective item. Lemma 5(b) in Appendix G gives that the event
$\mathscr{A}_{2}=\Big{\{}\widetilde{M}=(1+o(1))e^{-\nu}k\Delta\Big{\}}$ (179)
@@
Lastly, we deduce the high-probability behaviour of $N_{0}$. For $\mathcal{M}\subset s$, letting $W^{(\mathcal{M})}$ denote the number of tests in which some item from $\mathcal{M}$ is placed, Lemma 7 in Appendix G gives that $\mathbbm{E}W^{(s)}=(1-e^{-\nu}+o(1))n$, and that $W^{(s)}=\mathbbm{E}W^{(s)}+\delta^{\prime}n\sqrt{\frac{\ell}{k}}$ holds with probability at least $1-2\exp(-2\nu^{-1}\delta^{2}n)$ for any $\delta>0$ and for some $\delta^{\prime}<\delta$. Therefore, by letting $\delta=\Theta(\frac{1}{\sqrt{k}})$ we obtain $W^{(s)}=(1-e^{-\nu}+o(1))n$, which implies $N_{0}=n-W^{(s)}=ne^{-\nu}(1+o(1))$, holds with $1-o(1)$ probability. Thus, the event
$\mathscr{A}_{3}:=\big{\{}N_{0}=ne^{-\nu}(1+o(1))\big{\}}$ (180)
@@
$\mathcal{K}^{\prime}_{C,\zeta}=\Big{\{}j\in[k]:M_{j}^{\prime}=\frac{Cn\nu e^{-\nu}}{k},M_{j0}^{\prime}=\frac{\zeta\cdot Cn\nu e^{-\nu}}{k}\Big{\}},$ (181)
and let $k_{C,\zeta}^{\prime}=|\mathcal{K}^{\prime}_{C,\zeta}|$. To quantify the difference between $\mathcal{K}_{C,\zeta}$ and $\mathcal{K}^{\prime}_{C,\zeta}$, we provide the following observation
$\mathcal{K}^{\prime}_{C,\zeta}\subset\mathcal{K}_{C,\zeta}\cup\big{\{}j\in[k]:M_{j}^{\prime}<M_{j}\big{\}}.$ (182)
Combining with the event $\mathscr{A}_{1}$ in (177), this implies
$k^{\prime}_{C,\zeta}\leq k_{C,\zeta}+C_{0}\log k.$ (183)
Given $M_{j}^{\prime}=\frac{Cn\nu e^{-\nu}}{k}$, we have $M_{j0}^{\prime}\sim\text{Bin}(\frac{Cn\nu e^{-\nu}}{k},\rho)$ based on the randomness of $\mathbf{Z}$, so $M_{j0}^{\prime}=\frac{\zeta\cdot Cn\nu e^{-\nu}}{k}$ can be easily analyzed. Thus, the major challenge lies in analyzing the event $M_{j}^{\prime}=\frac{Cn\nu e^{-\nu}}{k}$, and we now proceed to study the distribution of $M_{j}^{\prime}$.
The distribution of $M_{j}^{\prime}$: We follow the idea in [16, Lemma 3.3] of conditioning on $\mathbf{T}_{s}$ (i.e., $\mathbf{T}_{[k]}$) and using the symmetry of the test design to deduce that the conditional assignments of tests to items are uniform, i.e., every partition of $\mathbf{T}_{s}$ into $k$ sets of size-$\Delta$ is equally likely. In particular, note that $\widetilde{M}$ defined in (178) is a known quantity from the given $\mathbf{T}_{s}$, and given $\widetilde{M}$, the number of such tests assigned to any given $j\in s$, which has been defined as $M_{j}^{\prime}$, follows a hypergeometric distribution: There are $\Delta$ draws (without replacement) from $k\Delta$ objects (that represent the overall $k\Delta$ placements in $\mathbf{T}_{s}$), $\widetilde{M}$ of which are of a special type (namely, the placement is to a test that is connected to precisely one defective item). For instance, regarding the allocation of these $k\Delta$ placements to item 1, we have $(M_{1}^{\prime}|\mathbf{T}_{s})\sim\mathrm{Hg}(k\Delta,\widetilde{M},\Delta)$. To understand the behavior of the first $k^{\zeta}$ items simultaneously, we interpret the above-mentioned uniform allocation from $\mathbf{T}_{s}$ as being done sequentially:
@@
  • Given $\mathbf{T}_{s}$ and $\mathbf{T}_{1}$ (the placements for item $1$), we allocate a random size-$\Delta$ subset of the reduced multi-set $\mathbf{T}_{s}\setminus\mathbf{T}_{1}$ to item $2$, which gives $M_{2}^{\prime}|(\mathbf{T}_{s},\mathbf{T}_{1})\sim\mathrm{Hg}\big{(}(k-1)\Delta,\widetilde{M}-M_{1}^{\prime},\Delta\big{)}$ (where $M^{\prime}_{1}$ is determined via $\mathbf{T}_{s},\mathbf{T}_{1}$).
@@
Similarly to the analysis of the Bernoulli design, we only consider this procedure up to index $k^{\xi}$, since going all the way up to $k$ would introduce non-negligible dependencies. For $\xi\in(\frac{1}{2},1)$ we have $k-j=(1+o(1))k$ for any $0\leq j\leq k^{\xi}$. Then, we utilize the event $\mathscr{A}_{2}$ in (179) that specifies $\widetilde{M}=(1+o(1))e^{-\nu}k\Delta$, and write
$(1+o(1))e^{-\nu}k\Delta=\widetilde{M}\geq\widetilde{M}-\sum_{j=1}^{i}M_{j}^{\prime}\geq\widetilde{M}-k^{\xi}\Delta=(1+o(1))e^{-\nu}k\Delta$ (184)

### 23 Oct 2024 v2

@@ Appendix D $-\ell$ Achievability Analysis for Bernoulli Designs
Bounding the overall failure probability for $\ell\in[1,\frac{k}{\log k}]$: We further take a union bound over $\ell\in[1,\frac{k}{\log k}]$. Suppose that the condition (D.67) holds for all $\ell\in[1,\frac{k}{\log k}]$, i.e.,
$n\geq(1+\eta_{1})\frac{k}{\nu e^{-\nu}}\max_{1\leq\ell\leq\frac{k}{\log k}}\max_{\begin{subarray}{c}0<C<C_{0}\\ 0<\zeta<1\end{subarray}}\min\Big{\{}\frac{\log\frac{k}{\ell}}{f_{1}(C,\zeta,\rho)},\frac{\log\frac{p}{\ell}}{\hat{f}_{2}(C,\zeta,\rho)}\Big{\}}$ (D.70)
$\iff$ $n\geq\frac{(1+\eta_{1})k\log\frac{p}{k}}{(1-\theta)\nu e^{-\nu}}\frac{1}{\min_{C\in(0,C_{0}),\zeta\in(0,1)}\max\{\frac{1}{\theta}f_{1}(C,\zeta,\rho),\hat{f}_{2}(C,\zeta,\rho)\}},$ (D.71)
@@
where in (D.71) we observe that the maximum over $\ell$ is attained at $\ell=1$ and use $\log p\sim\frac{1}{1-\theta}\log\frac{p}{k},\leavevmode\nobreak\ \log k\sim\frac{\theta}{1-\theta}\log\frac{p}{k}$. Then, we can bound the probability that the restricted MLE decoder fails as
$\mathbb{P}({\rm err})$ $\leq\mathbb{P}(\mathscr{A}_{1}^{c})+\sum_{\ell=1}^{k/\log k}\mathbb{P}\Big{(}\big{\{}\big{|}s\setminus\widehat{S}^{\prime}\big{|}=\ell\big{\}}\cap\mathscr{A}_{1}\Big{)}$ (D.72)
$\leq o(1)+\sum_{\ell=1}^{k/\log k}\Big{(}k^{-5}+O((\ell\log k)^{2})\hat{P}_{\ell}\Big{)}$ (D.73)
$\leq o(1)+\Big{[}\sum_{\ell=1}^{\log k}(\ell\log k)^{-3}+\sum_{\ell=\log k}^{k/\log k}k^{-3}\Big{]}=o(1),$ (D.74)
@@
where (D.72) follows from the union bound, (D.73) follows from (D.69), and (D.74) follows by substituting $\hat{P}_{\ell}$ (given in (D.64)). In conclusion, if (D.71) holds for some $\eta>0$, then the restricted MLE decoder fails to recover $s=[k]$ with $o(1)$ probability.
@@
Simplifying $\hat{f}_{2}(C,\zeta,\rho)$ to $f_{2}(C,\zeta,\rho)$: Recall that $\hat{f}_{2}(C,\zeta,\rho)$ is given in (D.57). The aim of this step is to show that the minimum over $(C,\zeta)$ in (D.71) is not attained in the domain of $(1-2\zeta)C<2\rho-1$, and thus we can safely simplify $\hat{f}_{2}(C,\zeta,\rho)$ to $f_{2}(C,\zeta,\rho)$ in the threshold (D.71). Consider $(C,\zeta)$ satisfying $(1-2\zeta)C\leq 2\rho-1$, i.e.,
$\zeta\geq\zeta^{\prime}:=\frac{1}{2}+\frac{1-2\rho}{2C}\geq\frac{1}{2},$ (D.75)
@@
in which we have $\hat{f}_{2}(C,\zeta,\rho)=0$ (see (D.57) and recall that $\hat{f}_{2}(C,\zeta,\rho)$ is continuous), and that $f_{1}(C,\zeta,\rho)$ is monotonically increasing with respect to $\zeta$ in $[\frac{1}{2},1)$ (see (D.13)). Therefore, we can compare the values of $f_{1}(C,\zeta,\rho)$ and $f_{2}(C,\zeta,\rho)$ over $(C,\zeta)$ and $(C,\zeta^{\prime})$ as
$f_{1}(C,\zeta,\rho)\geq f_{1}(C,\zeta^{\prime},\rho),\leavevmode\nobreak\ \hat{f}_{2}(C,\zeta,\rho)=\hat{f}_{2}(C,\zeta^{\prime},\rho)=0,$ (D.76)
@@
$\max\Big{\{}\frac{1}{\theta}f_{1}(C,\zeta,\rho),\hat{f}_{2}(C,\zeta,\rho)\Big{\}}\geq\max\Big{\{}\frac{1}{\theta}f_{1}(C,\zeta^{\prime},\rho),\hat{f}_{2}(C,\zeta^{\prime},\rho)\Big{\}}.$ (D.77)
@@
$\min_{C\in(0,C_{0}),\zeta\in(0,1)}\max\Big{\{}\frac{1}{\theta}f_{1}(C,\zeta,\rho),\hat{f}_{2}(C,\zeta,\rho)\Big{\}}$ (D.78)
=\,= $\min_{\begin{subarray}{c}C\in(0,C_{0}),\zeta\in(0,1)\\ \zeta\leq\frac{1}{2}+\frac{1-2\rho}{2C}\end{subarray}}\max\Big{\{}\frac{1}{\theta}f_{1}(C,\zeta,\rho),\hat{f}_{2}(C,\zeta,\rho)\Big{\}}$ (D.79)
=\,= $\min_{\begin{subarray}{c}C\in(0,C_{0}),\zeta\in(0,1)\\ (1-2\zeta)C\geq 2\rho-1\end{subarray}}\max\Big{\{}\frac{1}{\theta}f_{1}(C,\zeta,\rho),f_{2}(C,\zeta,\rho)\Big{\}}$ (D.80)
@@
Therefore, to ensure (D.71), it suffices to have
$n\geq\frac{(1+\eta)k\log\frac{p}{k}}{(1-\theta)\nu e^{-\nu}}\frac{1}{\min_{C>0,\zeta\in(0,1)}\max\{\frac{1}{\theta}f_{1}(C,\zeta,\rho),f_{2}(C,\zeta,\rho)\}}$ (D.82)
for some $\eta>0$. This establishes the second term in (1.4).
@@ Appendix E $-\ell$ Converse Analysis for Near-Constant Weight Designs
In the near-constant weight design, each item is independently and uniformly placed at $\Delta=\frac{\nu n}{k}$ tests with replacement. As with the Bernoulli design, it suffices to analyze the MLE decoder (B.1) that finds $\widehat{S}$ having the most correct test outcomes (i.e., outcomes that would be obtained if there were no noise). By the symmetry of the test design with respect to re-ordering items, we can again consider $s=[k]$ as the underlying defective set without loss of generality.
@@
For $j\in s$ and $j^{\prime}\in[p]\setminus s$ we will use the notation $\mathcal{M}_{j},\mathcal{M}_{j0},\mathcal{M}_{j1},\mathcal{N}_{0},\mathcal{N}_{00},\mathcal{N}_{01}$ and the corresponding cardinalities (e.g., $N_{0}=|\mathcal{N}_{0}|$, $M_{j}=|\mathcal{M}_{j}|$) from Corollary B.1. These sets and their cardinalities are deterministic given $\mathbf{X}_{s}$ and $\mathbf{Y}$. We consider $\ell=1$, and say that $(C,\zeta)\in(0,\infty)\times(0,1)$ is feasible if $\frac{Cn\nu e^{-\nu}}{k}$ and $\frac{\zeta\cdot Cn\nu e^{-\nu}}{k}$ are integers. Given a feasible pair $(C,\zeta)$, we will consider $\mathcal{K}_{C,\zeta}=\{j\in s:M_{j}=\frac{Cn\nu e^{-\nu}}{k},\leavevmode\nobreak\ M_{j0}=\frac{\zeta\cdot Cn\nu e^{-\nu}}{k}\}$ with cardinality $k_{C,\zeta}=|\mathcal{K}_{C,\zeta}|$ as in (B.7). For given $(j,j^{\prime})\in\mathcal{K}_{C,\zeta}\times([p]\setminus s)$, recall that $G_{j,j^{\prime},1}$ denotes the number of tests in $\mathcal{N}_{01}\cup\mathcal{M}_{j1}$ that contain the non-defective item $j^{\prime}$, and $G_{j,j^{\prime},2}$ denotes number of tests in $\mathcal{N}_{00}\cup\mathcal{M}_{j0}$ that contain the non-defective item $j^{\prime}$; see Corollary B.1.
We also introduce some additional definitions for analyzing the near-constant weight design. For $j\in s$ and $j^{\prime}\in[p]\setminus s$, we let $\mathcal{M}_{j,j^{\prime}}$ index the tests in $\mathcal{M}_{j}$ that contain item $j^{\prime}$, and $\mathcal{N}_{0,j^{\prime}}$ index the tests in $\mathcal{N}_{0}$ that contain item $j^{\prime}$. Then we note the equivalent interpretations of $G_{j,j^{\prime},1}$ and $G_{j,j^{\prime},2}$: $G_{j,j^{\prime},1}$ represents the number of positive tests in $\mathcal{N}_{0,j^{\prime}}\cup\mathcal{M}_{j,j^{\prime}}$, and $G_{j,j^{\prime},2}$ represents the number of negative tests in $\mathcal{N}_{0,j^{\prime}}\cup\mathcal{M}_{j,j^{\prime}}$.
In addition, for $j\in s$, we let $\mathcal{M}_{j}^{\prime}$ index the tests where item $j$ not only appears as the only defective but is also placed precisely once, and let its cardinality be $M_{j}^{\prime}=|\mathcal{M}_{j}^{\prime}|$. Compared to Bernoulli designs, this is a new ingredient for handling near-constant weight designs whose random placements are done with replacement. Evidently, it holds that $M_{j}^{\prime}\leq M_{j}$, and $M_{j}^{\prime}<M_{j}$ occurs when item $j$ is placed in some test more than once but is still the only defective. This situation only occurs for a small number of defective items, as we will formalize via event $\mathscr{A}_{1}$ below.
@@
Similarly to the converse analysis for Bernoulli design, by Corollary B.1, MLE fails if for some feasible $(C,\zeta)\in(0,e^{\nu})\times(0,1)$, the following two conditions (C1) and (C2) simultaneously hold:
@@
  • (C2) There exist $j\in\mathcal{K}_{C,\zeta}$ and $j^{\prime}\in\{k+1,\dotsc,p\}$ such that
    $G_{j,j^{\prime},1}-G_{j,j^{\prime},2}>(1-2\zeta)\frac{Cn\nu e^{-\nu}}{k}.$ (E.1)
@@
Before proceeding, we first construct three high-probability events, deferring the proofs that they have high probability to Appendix G. By Lemma G.1(a) in Appendix G, for some sufficiently large constant $C_{0}>0$, the event
$\mathscr{A}_{1}=\Big{\{}|\{j\in[k]:M_{j}^{\prime}<M_{j}\}|\leq C_{0}\log k\Big{\}}.$ (E.2)
@@
$\widetilde{M}:=\sum_{j=1}^{k}M_{j}^{\prime},$ (E.3)
which represents the total number of tests having exactly one defective placement (and no repeated placements of that defective item). Lemma G.1(b) in Appendix G gives that the event
$\mathscr{A}_{2}=\Big{\{}\widetilde{M}=(1+o(1))e^{-\nu}k\Delta\Big{\}}$ (E.4)
@@
Lastly, we deduce the high-probability behavior of $N_{0}$. For $\mathcal{M}\subset s$, letting $W^{(\mathcal{M})}$ denote the number of tests in which some item from $\mathcal{M}$ is placed, Lemma G.3 in Appendix G gives that $\mathbbm{E}W^{(s)}=(1-e^{-\nu}+o(1))n$, and that for any $\delta>0$, it holds with probability at least $1-2\exp(-2\nu^{-1}\delta^{2}n)$ that $W^{(s)}=\mathbbm{E}W^{(s)}+\delta^{\prime}n\sqrt{\frac{\ell}{k}}$ for some $|\delta^{\prime}|<\delta$. Therefore, by letting $\delta=\Theta(\frac{1}{\sqrt{k}})$, we obtain $W^{(s)}=(1-e^{-\nu}+o(1))n$, which implies $N_{0}=n-W^{(s)}=ne^{-\nu}(1+o(1))$, holds with $1-o(1)$ probability. Thus, the event
$\mathscr{A}_{3}:=\big{\{}N_{0}=ne^{-\nu}(1+o(1))\big{\}}$ (E.5)
@@
$\mathcal{K}^{\prime}_{C,\zeta}=\Big{\{}j\in[k]:M_{j}^{\prime}=\frac{Cn\nu e^{-\nu}}{k},M_{j0}^{\prime}=\frac{\zeta\cdot Cn\nu e^{-\nu}}{k}\Big{\}},$ (E.6)
and let $k_{C,\zeta}^{\prime}=|\mathcal{K}^{\prime}_{C,\zeta}|$. To quantify the difference between $\mathcal{K}_{C,\zeta}$ and $\mathcal{K}^{\prime}_{C,\zeta}$, we observe that
$\mathcal{K}^{\prime}_{C,\zeta}\subset\mathcal{K}_{C,\zeta}\cup\big{\{}j\in[k]:M_{j}^{\prime}<M_{j}\big{\}}.$ (E.7)
Combining with the event $\mathscr{A}_{1}$ in (E.2), this implies
$k^{\prime}_{C,\zeta}\leq k_{C,\zeta}+C_{0}\log k.$ (E.8)
Given $M_{j}^{\prime}=\frac{Cn\nu e^{-\nu}}{k}$, we have $M_{j0}^{\prime}\sim\text{Bin}(\frac{Cn\nu e^{-\nu}}{k},\rho)$ based on the randomness of $\mathbf{Z}$, so the event $M_{j0}^{\prime}=\frac{\zeta\cdot Cn\nu e^{-\nu}}{k}$ can be easily analyzed. The more significant challenge lies in analyzing the event $M_{j}^{\prime}=\frac{Cn\nu e^{-\nu}}{k}$, and we now proceed to study the distribution of $M_{j}^{\prime}$.
The distribution of $M_{j}^{\prime}$: We follow the idea in [15, Lemma 3.3] of conditioning on $\mathbf{T}_{s}$ (i.e., $\mathbf{T}_{[k]}$) and using the symmetry of the test design to deduce that the conditional assignments of tests to items are uniform, i.e., every partition of $\mathbf{T}_{s}$ into $k$ sets of size-$\Delta$ is equally likely. In particular, note that $\widetilde{M}$ defined in (E.3) is a known quantity from the given $\mathbf{T}_{s}$, and given $\widetilde{M}$, the number of such tests assigned to any given $j\in s$, which has been defined as $M_{j}^{\prime}$, follows a hypergeometric distribution: There are $\Delta$ draws (without replacement) from $k\Delta$ objects (that represent the overall $k\Delta$ placements in $\mathbf{T}_{s}$), $\widetilde{M}$ of which are of a special type (namely, the placement is to a test that is connected to precisely one defective item). For instance, regarding the allocation of these $k\Delta$ placements to item 1, we have $(M_{1}^{\prime}|\mathbf{T}_{s})\sim\mathrm{Hg}(k\Delta,\widetilde{M},\Delta)$. To understand the behavior of the first $k^{\zeta}$ items simultaneously, we interpret the above-mentioned uniform allocation from $\mathbf{T}_{s}$ as being done sequentially:
@@
  • Given $\mathbf{T}_{s}$ and $\mathbf{T}_{1}$ (the placements for item $1$), we allocate a random size-$\Delta$ subset of the reduced multi-set $\mathbf{T}_{s}\setminus\mathbf{T}_{1}$ to item $2$, which gives $M_{2}^{\prime}|(\mathbf{T}_{s},\mathbf{T}_{1})\sim\mathrm{Hg}\big{(}(k-1)\Delta,\widetilde{M}-M_{1}^{\prime},\Delta\big{)}$ (where $M^{\prime}_{1}$ is determined via $\mathbf{T}_{s}$ and $\mathbf{T}_{1}$).
@@
Similarly to the analysis of the Bernoulli design, we only consider this procedure up to index $k^{\xi}$, since going all the way up to $k$ would introduce non-negligible dependencies. For $\xi\in(\frac{1}{2},1)$ we have $k-j=(1+o(1))k$ for any $0\leq j\leq k^{\xi}$. Then, we utilize the event $\mathscr{A}_{2}$ in (E.4) that specifies $\widetilde{M}=(1+o(1))e^{-\nu}k\Delta$, and write
$(1+o(1))e^{-\nu}k\Delta=\widetilde{M}\geq\widetilde{M}-\sum_{j=1}^{i}M_{j}^{\prime}\geq\widetilde{M}-k^{\xi}\Delta=(1+o(1))e^{-\nu}k\Delta$ (E.9)